Blossom, Inc., sells two types of water pitchers, plastic and glass. Plastic pitchers cost the company $16 and are sold for $31. Glass pitchers cost $25 and are sold for $46. All other costs are fixed at $982,800 per year. Current sales plans call for 14,000 plastic pitchers and 42,000 glass pitchers to be sold in the coming year.

How many pitchers of each type must be sold to break even in the coming year? (Use contribution margin per unit to calculate breakeven units.)
Plastic Pitchers

Glass Pitchers

Blossom, Inc., has just received a sales catalog from a new supplier that is offering plastic pitchers for $14. What would be the new contribution margin per unit if managers switched to the new supplier?

Plastic Pitchers Glass Pitchers

Contribution Margin Per Unit $ $
Answer #1

Weighted average contribution margin per unit = (31-16)*14000/56000+(46-25)*42000/56000 = 19.50

Break even unit = 982800/19.50 = 50400 Units

Plastic Pitchers = 50400*14/56 = 12600

Glass pitches = 50400*42/56 = 37800

Contribution margin per unit

Plastic pitchers Glass pitchers
Contribution margin per unit 31-14 = 17 46-25 = 21